OpenStudy (anonymous):

Frank can type a report in 7hr. James can type a report in 6hr. How long will it take if they are working together?

OpenStudy (anonymous):

guess!

OpenStudy (anonymous):

Trick question, longer, because sharing a keyboard will make things difficult.

OpenStudy (anonymous):

\[\frac{6\times 7}{6+7}\]

OpenStudy (anonymous):

42/13

OpenStudy (anonymous):

one the the job in m hours, other does the job in n hours. together then can do the job in \[\frac{m\times n}{m+n}\] hours

OpenStudy (anonymous):

yup, \[\frac{42}{13}\]

OpenStudy (anonymous):

Do I just reduce that

OpenStudy (anonymous):

it does not reduce because 13 is prime. you could write as a mixed number if you like

OpenStudy (anonymous):

3&3/13

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw hope method is clear

OpenStudy (anonymous):

I am starting to get it!

OpenStudy (anonymous):

good. i could write an explanation of you like, but the procedure is simple

OpenStudy (anonymous):

I have trouble setting up the problem!

OpenStudy (anonymous):

An expericened accountant can prepare a certain complexity of a tax return in 11hr. Another can do it in 22hr. How long will it take with them working together?

OpenStudy (anonymous):

\[\frac{11\times 22}{11+22}\]

OpenStudy (anonymous):

242/33

OpenStudy (anonymous):

now it is just a computation

OpenStudy (anonymous):

so basically just mutiply on top and add on bottom?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then divide 33by 242?

OpenStudy (anonymous):

one has a rate of \[\frac{1}{11}\] other has a rate of \[\frac{1}{22}\] combined rate is \[\frac{1}{11}+\frac{1}{22}=\frac{22+11}{22\times 11}\]

OpenStudy (anonymous):

you want to solve rate times time = 1 (one job) so time is reciprocal of combined rate which is \[\frac{22\times 11}{22+11}\]

OpenStudy (anonymous):

this reduces a lot

OpenStudy (anonymous):

get \[\frac{22}{3}\]

OpenStudy (anonymous):

yes I got that to

OpenStudy (anonymous):

or \[7\tfrac{1}{3}\] if you prefer

OpenStudy (anonymous):

or in time , 7 hours and twenty minutes

OpenStudy (anonymous):

Thank you so much I understand it now!